OWG Meeting 18-Jun-2014

From GlueXWiki
Jump to: navigation, search

Location and Time

Room: CC F326-327

Time: 1:30pm-2:30pm


(if problems, call phone in conference room: 757-269-6460)

  1. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 120390084.
  2. To join via a Web Browser, go to the page https://bluejeans.com/120390084.
  3. To join via phone, use one of the following numbers and the Conference ID: 120390084
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. More information on connecting to bluejeans is available.

*note: you'll be muted when first connected to the meeting. press *4 to unmute


  1. Announcements
  2. SQLite DB + EPICS translation table (Orlando)
  3. Major systems reports
    • Trigger
    • DAQ
    • Monitoring
    • Controls
  4. Mantis Task Tracker

Previous Meeting



David L. (chair), Curtis M., Sean D., Vardan G., Sergey F., Alex S., Hovannes E., Orlando S., Beni Z., Simon T., Bryan M., Mark D., Dave A.

SQLite DB + EPICS translation table

Orlando presented some work he has done to create scripts capable of merging two SQLite database files containing translation table information. The definitive source for translation table information is in an SQLite file maintained in the subversion repository. This may be copied and modified simultaneously by different detector groups. The modifications and additions therefore must be merged back into a single DB. The merge is non-trivial due to the structure of the tables and use of primary keys. Orlando showed some examples of how to use the scripts and agreed to put them in the subversion repository so others may access them.

n.b. Since the meeting, Hovanes has added these to the repository here:


Major subsystems reports


  • The connection of the CAEN 1495 external pulser to the FCAL and TS has been finalized. This is used to trigger pulser LEDs
  • Work is shifting to the TOF which is the higher priority now. Alex S. is working with the FSU folks to get a working DAQ that self triggers from the TOF itself


  • A major rework of the L1 trigger and front-end module configuration file format is nearly complete and will be implemented next week.
  • Sergey F. and David have scheduled a weekly timeslot (Thursday mornings 9am-11am) to due large scale DAQ system testing.
    • The first LSD test revealed an issue when trying to launch 50 crates. Several (roughly half) would not complete connection to the platform. This may be due to multiple NICs
    • Vardan has generated a single threaded version of the relevant code that may solve much of the problem.
  • Sergey F. has started looking into the insertion of non-DAQ, "user" events into the data stream.


  • David has been working on a Python module for parsing the COOL configuration. This can be used for things like extracting the ET connection parameters for the ER-made ET system.
  • Work has also been started on modifying the EVIO parsing code to handle event stacking by the ET system. This can happen when reading from the ET, but not from files.
    • Dave A. will talk to Carl about whether they really want CODA to pass us stacked events or not.


  • Hovanes and Vanik have successfully tested the controls for the collimator table motion
  • Vanik has finished the hodoscope voltage controls
  • Vanik and Nerses have refactored a number of FDC variable names
  • Vanik has made modifications to the FDC frame to indicate alarms
  • Orlando has put the discriminator channel mapping into tt.db
  • Hovanes has been working on some reorganization of some of the GUIs.

Action Items